which number is not a solution to the equation
(2x-3)(x-11)(x/3+7)=0?
A.-21
B.11
C.-3/-2
D.3/2

Found 2 solutions by Alan3354, Theo:
Answer by Alan3354(69443)
Answer by Theo(13342)

if a * b * c = 0, then either a can be equal to 0 or b can be equal to 0 or c can be equal to 0 and the equation will be true because you will get 0 = 0.

take each factor in turn and set it equal to 0 and solve for the unknown variable.

2x-3 = 0
add 3 to both sides to get:
2x = 3
divide both sides by 2 to get
x = 3/2

that's solution D.

x-11 = 0
add 11 to both sides.
x = 11

that's solution B.

x/3 + 7 = 0
subtract 7 from both sides to get:
x/3 = -7
multiply both sides by 3 to get:
x = 21

that's solution A.

Look's like the solution that is not a solution to your problem is C.